OpenStudy (anonymous):

Two investments were made totaling $15,000. For a certain year these investments yeild $1326 in simple interest. Part of the 15,000 was invested at 14% and part at 16%. Find the amount invested at each rate

OpenStudy (mathstudent55):

For simple interest, the interest earned, i, is i = prt where p = principal, r = rate, t = time

OpenStudy (mathstudent55):

Let x = amount invested at 14%, and y = amount invested at 16%. Then using i = prt, you get: For the amount invested at 14% \(i_{14} = x(14\%)(1) = 0.14x\) For the amount invested at 16% \(i_{16} = y(16\%)(1) = 0.16y \) The total interest earned is \(i_{14} + i_{16}\) and equals $1326. 0.14x + 0.16y = 1326 Eq. 1 The total amount invested is x + y = 15000 Eq. 2 Solve equations 1 and 2 as a system of equations to find the amounts.

OpenStudy (anonymous):

Ahh I see.. Alright I'll see what I've got from it.

OpenStudy (anonymous):

ok thanks for the help ^^